FEL Upgrade Project Weekly Brief

March 18-22, 2002




We are pleased to announce that both ONR and DOE have signed the new Memorandum of Agreement (MOA) for the FY02/03 activities on the 10 kW upgrade and the related FEL technology development grants from the Joint Technology Office.

We are also pleased to acknowledge the efforts of our colleague Michael Kelley, who was notified this week of a winning grant from the Defense University Instrumentation Program (DURIP). The $467k grant will cover the development of pulse compression and stretching equipment for the FEL user labs.


The project monthly reports for February 2002 were distributed to the DOE, ONR and AFRL program offices.

Discussions were held this week with the ONR contract monitor on a draft agenda for the upcoming May 1-3, 2002, FEL project semiannual review. We plan to distribute the agenda next week for comments by the review committee.

F. Dylla and the JLab Director, Christoph Leemann, attended the quarterly meeting of Virginia’s Research and Technology Advisory Committee (VRTAC) on March 18-19th, which hosted a briefing for the Virginia Congressional delegation on R&D priorities for the Commonwealth VRTAC continues to support Jefferson Lab’s nuclear physics and FEL programs.

WBS 3 (Beam Physics):

During discussions this week of the vacuum chamber design we learned that the year-old beam stay clears could be modified with no significant impact at this point to extend performance to maintain capability for the 100 kW level. This situation was addressed with our usual methods of give-and-take between the chief lattice design and the hardware designers and brought to rapid and satisfactory conclusion.

Having agreed on a means of dealing with issues of tracking and matching amongst endloop dipoles, we began discussions with WBS 9(George B. & co.) and AES to have it happen. We also started writing it up to document the rather arcane UV geometry and allow people to review our conclusions.



WBS 4 (Injector):

Continued polishing the ball cathode and support tube. Ordered ceramic tubes for the conditioning resistor, design ready for detailed piece drawings. Design for the shielded bellows is nearly complete.

Gun HVPS Tank - The tank is being manufactured by Industrial Alloy Fabrication of Richmond. It is due 10 May. Parts have been ordered and are being received for the connections into the tank. The mechanism that combines the conditioning and running resistors has been designed and is being detailed. The large ceramic cylinder is on order.

PARMELA simulations of upgrade gun performance show an improvement in the longitudinal and transverse emittances after fine tuning of the buncher gradient and solenoid field. The results up to the end of the 1/4 Cryounit are: z-emittance ~18 deg-keV, x-emittance ~5.7 mm-mrad, and y-emittance ~4.5 mm-mrad. Work is in progress towards transporting the beam to the end of the LINAC.

WBS 6 (RF):

Quarter HVPS - The wiring and documentation for the Upgrade continues. Checkout should be able to start 1 April.

Quarter Klystrons - The drawings and many of the fittings have been provided to the Machine Shop for the 2 Klystron Carts they are building. Progress on the LCW piping modifications has stopped until the valve issue can be resolved and the welders are available after the CEBAF shutdown.

WBS 8 (Instrumentation):

A big thank you to the operations group for the additional help during the CEBAF shutdown. The operators are assisting in the BPM detector testing, assembly and testing of beam viewer cameras, and fabrication of a number of VME boards required for the upgrade.

The final push is on to complete all tasks associated with beam viewer system. This effort will complete mechanical assembly, camera rework, control chassis fab and pre assembly of camera support structure. Upon completion, the remaining viewer assemblies will be delivered to the magnet girder assembly area in the Test Lab. Additionally the CCD elements which suffered radiation damage are being replaced.

All of the 3" Beam Position Monitors (BPM) are being retested prior to delivery to the Test Lab, of the 44 detectors (176 feedthrus) 2 feedthrus were found to be intermittent. These will have to be reworked prior to installation. The test was to prevent any problem detectors from being installed.

The new Beam Loss Monitor (BLM) specifications were reviewed again. There is a need for a "diagnostic" mode for the BLMs but it was decided that installing an additional 6 to 8 "floaters", which are not interlocked to the Fast Shut-Down (FSD) system, provides the best solution. By having independent control over the tube voltage one can get 6 decades of "gain", this can not be accomplished on the interlocked tubes without compromising the system integrity.

The DC power trim channels have been assigned and are being inputted in the web site for generation of cable labels. The additional trim rack should be installed in the gallery in the next few weeks. A meeting was also held regarding the "super trim" power supply, these are 75Volt/20Amp bipolar supplies. Fourteen (14) of these are required for the IR/UV machine, we will order two (2) spares. The only remaining procurement is for seven (7) 100 amp power supplies.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• At Magnet Enterprises International (MEI) in Oakland CA. They are preparing the third coil of for potting and winding the fourth coil. With success of their technique, they are making one more winding form to up the production rate to be potting limited.

• I inspected the first two coils and found that they are essentially acceptable, having passed hi pot test. Some dressing of some high spots in the resin clamped area and some fill in with epoxy to some superficial dry areas was authorized.

• Return legs are being prepared for final machining at a very competent shop. Top and bottom slabs are at the heat treat facility. Field clamps remain in process. Covers look beautiful.

• MEI is attempting to have one complete magnet by the review.

Injector Dipoles (DU/DV)

• Master Machine continues manufacturing.

• WANG NMR successfully wound the wound GV Coil and made the GU winding tool.

Arc Dipoles (GY, GX, GQ)

• We completed discussions with offerors and requested "Best and Final Offers" that are due Monday.

• Wang NMR of CA continues making the GY coil winding mandrel and designing the potting mold for the GY coils. They have wound all the copper onto reels to allow rapid insulation and winding for these four-in-hand coils.

• Bosma Machine has the GY core pieces in final machining. They are final machining the return legs.


3 inch quad (QX)

• Production measurements continued on the remaining QX Quadrupoles. We are concentrating on the "skinny" quads that David Douglas wants to assign to the regions between cryomodules.

Trim Quad (QT)

• Milhous Control of Virginia reports that a first article has been delayed to next week.

• The wire manufacturer re spooled the wire in an untwisted way and it is at Milhous.

Sextupole (SF)

• DULY Research is putting together the report on their successful design.

Octupole (OT)

• Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

• In checking the chamber configuration, we found that the chamber passes the 10 kW beam easily but restricts the beam in two places for 100 kW in the second arc. We will change the design to a wider configuration in those positions. Design at AES continues, drawing up the remainder of the arc chambers.

• Design continued on the details for the region after the cryomodules as well as determining the exact position of the cryomodules

• LCW system design continues. This week we concentrated on fabrication practice and design for commissioning. Since the CEBAF System is already pure, we have to minimize disruption of its quality as we open our system to it.

WBS 10 (Wiggler):

Degaussing the dispersion section changed the remanent fields but not the fields at full currents. The field integral was unchanged. We spent the week looking for possible causes of the mysterious field integral change. Despite changing many things and disassembling and reassembling the magnet several times, the field integral remained unchanged. We still have no explanation for the change from last years measurements. Since the field integral seems quite stable right now we will just do the final measurements with no more changes. The final result of the increased field integral is that the correction coil will have no headroom when operating the dispersion section at full strength. Since the wiggler coils have headroom this is not a big problem.

WBS 11 (Optics):

The UHV stepper motors (to be used in the optical cavity and collimator) began testing this week. The I & C group were able to create settings for their stepper motor controller that produced smooth operation. We determined (to an accuracy of +/- 60 nm) that there is no backlash when reversing direction. Tests of the piezo actuator will be performed today, using a power supply. We had hoped to use the amplifiers purchased from the vendor, but the power supply failed and it is back at the factory for repair. We anticipate that it will be on-site next week.

Design efforts are centered around completing the optical cavity vacuum chambers, as well as some of the internal mounting hardware, the details of the insertable mirror holder, and the turning mirror cassettes.

The procurement of the ultrafast laser was switched from sole source to commercial bid.

Bids are due back early next week, and we will then make an award.

We provided the Injector Source group with advice on how to improve the modelocking of their drive laser for the G0 project. They also used our optical inspection microscopy to look at one of their ball cathode assemblies.